OpenStudy (nino_jim):

you are meeting a friend at 5pm and she lives 240 miles away. the trip can be traveled entirely on a highway with a speed limit of 60mph. what time do you need to leave your house

OpenStudy (ffasinger):

do you know the formula?

OpenStudy (nino_jim):

no sorry

OpenStudy (ffasinger):

i am pretty sure it is distance dived by time . right @mathmale

OpenStudy (nino_jim):

s=d/t

OpenStudy (ffasinger):

i belive so..

OpenStudy (nino_jim):

yeah thats it

OpenStudy (mathmale):

You want to know how long it will take to drive 240 miles at 60 mph. How would you find that? Again , review the formula d = r t, where d=distance, r=rate (in mph), and t=time (in hrs).

OpenStudy (ffasinger):

thank you @mathmale

OpenStudy (mathmale):

Sure.

OpenStudy (mathmale):

Once you have the # of hours the trip will take, subtract that figure from 5:00 p.m. to find the departure time.

OpenStudy (nino_jim):

would it be 4:20

OpenStudy (ffasinger):

yup

OpenStudy (nino_jim):

you sure

OpenStudy (ffasinger):

yup

OpenStudy (nino_jim):

thanks so much @FFASinger

OpenStudy (ffasinger):

your welcome

OpenStudy (mathmale):

Please wait!! Can you actually drive 240 miles in 40 minutes? I can't. Ask yourself: does this result make sense? I think not. How long does it take to drive 240 mi at 60 mph?

OpenStudy (ffasinger):

oh id did that wrong !! oops